1. Lucas is the worlds WORST sno cone maker. He drops 78% of the snow cones he makes. Its a terrible waste of sno. Anyway,
Answer #1

n=20 (Binomic distribution a 1. p=0.78 (781) pin) =CBM (1-b41-9> 20,10-78) (02*; =0,1,2, --1re. $(7= 12) = 20 (2.78² (0.28)